Two parallel straight wires 10.0 cm apart carry currents in opposite directions (Fig. 20โ€“24). Current I_1=5.0 A is out of the page, and I_2=7.0 A is into the page. Determine the magnitude and direction of the magnetic field halfway between the two wires.

1
Expert's answer
2022-03-24T16:36:46-0400

Net magnetic field

"B=B_1+B_2"

"B=\frac{\mu _0}{2\pi}(\frac{i_1}{r_1}+\frac{i_2}{r_2})"

"B=\frac{4\pi}{2\pi}(\frac{5}{0.05}+\frac{7}{0.05})"

"B=2\times(100+140)=480T"

Direction into the page